There are 10,000 sacks of rice. How many sacks of rice must be distributed to town A, B, and C if Town A has 15,000 families, Town B has 12,000, and Town C has 7,000?


1
Expert's answer
2021-03-01T15:57:53-0500

Total stacks of rice "n=10,000"

Total no. of famalies "=15,000+12,000+7,000=34,000"


Probability of town A families "P_1=\dfrac{15000}{34000}=0.44117"


Probability of town B families "P_2=\dfrac{12000}{34000}=0.3529"41


Probability of town C families "P_3=\dfrac{7000}{34000}=0.20588"



Number of stacks distributed to town "A= P_1\times n=0.44176\times 10000=4411.76"


Number of stacks distributed to town "B= P_2\times n=0.352941\times 10000=3529.41"


Number of stacks distributed to town "C= P_3\times n=0.20588\times 10000=2058.8"